# MalPurifier: Enhancing Android Malware Detection with Adversarial Purification against Evasion Attacks

### v1

@@ Abstract
Machine learning (ML) has gained significant adoption in Android malware detection to address the escalating threats posed by the rapid proliferation of malware attacks. However, recent studies have revealed the inherent vulnerabilities of ML-based detection systems to evasion attacks. While efforts have been made to address this critical issue, many of the existing defensive methods encounter challenges such as lower effectiveness or reduced generalization capabilities. In this paper, we introduce a novel Android malware detection method, MalPurifier, which exploits adversarial purification to eliminate perturbations independently, resulting in attack mitigation in a light and flexible way. Specifically, MalPurifier employs a Denoising AutoEncoder (DAE)-based purification model to preprocess input samples, removing potential perturbations from them and then leading to correct classification. To enhance defense effectiveness, we propose a diversified adversarial perturbation mechanism that strengthens the purification model against different manipulations from various evasion attacks. We also incorporate randomized “protective noises” onto benign samples to prevent excessive purification. Furthermore, we customize a loss function for improving the DAE model, combining reconstruction loss and prediction loss, to enhance feature representation learning, resulting in accurate reconstruction and classification. Experimental results on two Android malware datasets demonstrate that MalPurifier outperforms the state-of-the-art defenses, and it significantly strengthens the vulnerable malware detector against 37 evasion attacks, achieving accuracies over 90.91%. Notably, MalPurifier demonstrates easy scalability to other detectors, offering flexibility and robustness in its implementation.
@@ 1 Introduction
Unfortunately, ML-based detectors are vulnerable to adversarial examples, which are created by modifying non-functional instructions in executable programs of existing malware [8, 9, 10]. Adversarial examples can enable a range of attacks, including evasion attacks [11, 12, 13], poisoning attacks [14, 15, 16], or a combination of both [17]. In this study, we specifically narrow our focus to evasion attacks, which are designed to deceive ML-based detection during the testing phase.
@@
Another defense technique, known as adversarial purification [24, 25, 26], aims to remove potential perturbations from input samples, resulting in purified samples that can be correctly classified by the target classifier. The purification model is usually trained independently of the classification model and does not necessarily require class labels [27]. As a result, it can mitigate unseen threats in a plug-and-play manner without re-training the target classifier [28], leading to less training overhead and more flexible employment. Nevertheless, existing purification solutions for image classification are not easily applicable to Android malware detection due to significant differences. The input space of the image classifier is continuous, while malware features are often discrete [29]. Hence, it is more challenging to develop a robust Android malware detector.
Inspired by the complementary strengths of existing approaches in image classification, we propose a customized adversarial purification-based method to defeat evasion attacks in Android malware detection, as illustrated in Fig. 1. This novel mechanism has the following highlighted features that enable it to eliminate perturbations in adversarial malware and restore it closer to its original form: (i) We first generate diversified adversarial malware samples by adding perturbations in a progressively increasing manner to represent varying evasion attacks. This approach enables the purification model to fully acquire the capability to effectively purify these perturbed samples. (ii) We further injected “protective noises” into the benign samples during the training of the purification model. This step enhances the distinction between benign and adversarial samples, preventing excessive purification of benign samples. (iii) Based on these generated samples with their original versions, we then train a Denoising AutoEncoder (DAE) based purification model with a customized loss function in an unsupervised manner, allowing us to accurately pre-process the samples for classification purposes. (iv) More importantly, the proposed method is non-intrusive, computationally efficient, and easily scalable to different detection models. The main contributions of this paper can be summarized as follows:
  • Attack mitigation with adversarial purification. We propose a robust Android malware detection method, termed MalPurifier, that leverages a plug-and-play DAE model to purify adversarial malware. To the best of our knowledge, this is the first attempt to use adversarial purification to mitigate evasion attacks in the Android ecosystem.
@@
  • Trade-off between robustness and detection accuracy. Our proposed mechanism injects a diversified range of perturbations into malware samples, ranging from no injection to worst-case perturbations, to enhance robustness against different types of attacks. Moreover, we also implement safeguards to prevent false positives on benign samples during the purification process.
@@
  • Experimental validation on public datasets. We compare MalPurifier with the state-of-the-art (SOTA) methods via Drebin [30] and Androzoo [31] datasets. Experimental results show that MalPurifier significantly outperforms other defenses against attacks with less overhead required. Finally, we release the source code at https://github.com/SEU-ProactiveSecurity-Group/MalPurifier.
@@ 2.3 Adversarial Purification
The fundamental concept behind adversarial purification is to preprocess the input data directly, preventing any embedded adversarial components from feeding into the target model, so that the influence of attacks can be mitigated. These methods are widely regarded as model-agnostic and highly efficient, making them easy to train and utilize while demonstrating strong generalization capabilities.
@@ 3.1 Threat Model
To perform analysis on potential attacks and develop countermeasures, we consider the threat model in terms of explicit assumptions regarding the attacker’s capabilities and knowledge of the target system. Especially, we discuss prevailing attack scenarios specified by three different threat models in this section.
@@ 3.1.1 Black-Box Attacks
The attacker has no knowledge about the malware detector $f$ or the purifier $g$, but can only obtain the prediction results from the target system. Given a malware example $z$, we can have $x=\phi(z)$, and the attacker aims to modify it for producing wrong prediction such that
@@ 3.1.2 Grey-Box Attacks
The adversary has knowledge of the original malware classifier $f$ but is not aware of the purifier $g$ (i.e., oblivious attack). Hence, the adversary’s goal is only to fool the unsecured model $f$ using adversarial examples generated from malicious examples. Formally, given a malware example $z$ with its feature representation $x=\phi(z)$ and label $y=1$, the attacker will modify it to obtain the adversarial feature representation $x^{\prime}$ that can evade detection, by solving
@@ 3.1.3 White-Box Attacks
In this attack scenario, the adversary has full knowledge of the model $f$ and is aware of the structure of the purifier $g$ (i.e., adaptive attack). In other words, the adversary needs to mislead both the malware detector and adversarial purifier simultaneously by perturbing $x$ into $x^{\prime}$. It is worth noting that the sample will be first feed into the purifier and its output will be then classified by the malware detector.
Thereby, given a malware feature-label pair $(x,y)$ in this study, the attacker needs to solve
@@ 3.2 Defense Formulation
  • High-accuracy Android malware detection. Developing a malware detection model that can classify clean samples into benign or malicious with high accuracy, that is, $\varphi_{\theta}(x)=y$, in which $x$ can be the feature representation of a normal Android application or a malware sample, and thus $y=0$ or $1$. This model can be easily obtained by utilizing some ML algorithms (i.e., Deep Neural Network (DNN)), which have shown promising results with 99% accuracy in their laboratory settings [40, 12, 6].
@@ 4 The MalPurifier Approach
In this section, we provide an in-depth exploration of the MalPurifier approach. We begin by offering an overview that encompasses the processes of model training and inference. Subsequently, we delve into the mechanisms of diversified adversarial perturbation and protective noise injection, which are employed to generate training data for the purification model. Finally, we elaborate on the establishment of a label-independent DAE-based purification model, which plays a crucial role in producing purified samples and thereby contributing to accurate predictions.
@@
4.1 Defense Architecture
The MalPurifier plays two important roles, i.e., adversarial purification and malware detection. First, it performs input preprocessing in a label-independent manner and generates reconstructed representations in new forms. Especially, adversarial samples are restored to their original versions, while the characteristics of clean data remain unchanged. Second, it determines the benignity of the inputs by analyzing the prediction results of the DNN model on the processed samples. Fig. 2 depicts an overview of the MalPurifier architecture and explains how the functionally different components are jointly optimized.
As we can see, the proposed method consists of three main modules: (i) a feature extractor $\phi(\cdot)$ that maps an Android application into a feature vector, (ii) an adversarial purifier $g$ that processes samples via a DAE model $\psi_{\vartheta}$, and (iii) a malware detector $f$ that uses a DNN model $\varphi_{\theta}$ for detection. In the training phase, we initiate the process by extracting features from multiple batches of clean (or natural) examples, which include both benign and malicious examples without manipulation. These features are then fed into the DNN model, which iteratively updates its parameters to minimize the loss function. Once the DNN model is effectively trained, it exhibits exceptional classification accuracy when presented with clean inputs. At the same time, as a detection model, it often requires a substantial amount of training samples. To reduce costs, we will not retrain this model in this study, in contrast to adversarial training methods.
Different from the DNN, the DAE learns compressed representations of input data and reconstructs the original (clean) data from noisy versions. The performance of a DAE model is significantly influenced by the training data, especially the level and type of noise introduced during training. To enhance the DAE’s ability to handle various adversarial perturbations in input data, we propose a diversified adversarial perturbation mechanism, as demonstrated in Section 4.2. This mechanism generates a range of adversarial malware samples with varying attack strengths, thereby improving the model’s effectiveness in mitigating evasion attacks. Furthermore, in this study, we introduce a noise injection mechanism (see Section 4.3) to safeguard benign samples from excessive or erroneous purification by the DAE, hence preserving their original feature patterns. Finally, we incorporate reconstruction loss and prediction loss into the DAE model, and feed both generated adversarial samples and noisy benign samples, along with their corresponding clean forms into it. Note that we train the DAE model independent of the label space. More details of the DAE training can be found in Section 4.4.
In the phase of testing, the MalPurifier can be fed with benign, malicious and adversarial examples. By inputting an Android App into our method, one can obtain the prediction result (benign or malicious) as it passes through these modules. As depicted in the right part of Fig. 2, the MalPurifier demonstrates accurate identification of adversarial samples that may have evaded detection without purification.
@@ 4.2 Diversified Adversarial Perturbation
To train an adversarial purification model, we need to create perturbed malware samples first. Traditional adversarial training methods usually exploit strong white-box attacks to consider the target parameters $\theta$ and perturb the malware sample to create $x^{\prime}$, such that it will be misclassified by the detection model, i.e., $\varphi_{\theta}(x^{\prime})=0$. Since this method calculates perturbations using gradient directions specific to $\theta$ and stops manipulations when the classifier achieves misclassification, these perturbed malware samples cannot generalize well across different attack intensities.
To solve the aforementioned challenges, we design a diversified adversarial perturbation mechanism that can generalize across multiple kinds of attacks, thus enabling a robust defense approach. Specifically, this mechanism aims at increasing the feature space difference between a malware sample $x$ and its perturbed version $x^{\prime}$ instead of specific label information, which can be represented as
$\Delta(x,x^{\prime})=d(\mathcal{F}_{\theta}(x)|_{n},\mathcal{F}_{\theta}(x^{\prime})|_{n}),\ \ \mathrm{s.t.}x^{\prime}\in[\check{u},\hat{u}],$ (9)
where $\mathcal{F}_{\theta}(x)|_{n}$ denotes the internal representation of input data $x$ from the $n$th layer of the ML-based malware detector $f$, and $d(\cdot)$ quantifies the difference between the features of the original and perturbed samples. In this study, we employ Mean Square Error (MSE) as the distance metric. Furthermore, instead of directly maximizing this gap under constrained conditions, we add perturbations in a gradually increasing manner during the processing of input samples. This approach allows the generated samples to encompass a wider range of attack intensities, ensuring diversified and highly generalizable adversarial examples.
Algorithm 1 demonstrates the generation of diversified adversarial examples as the following steps. (i) We first sample a malware batch $(x_{i},y_{i}=1)_{i=1}^{N}$ in Line 2. (ii) Then, Line 3 sets the adversarial depth to a linear relationship with the batch index. (iii) In line 5, we do not inject any perturbations for the first batch of malware samples. (iv) For the other batch, we randomly transform $x_{i}$ as the initial point in Line 7. We derive the gradients $g_{t}$ in Line 10 according to the difference calculated in Line 9, thereby creating adversarial samples according to the preset depth in Line 11. To keep the generated adversarial samples practical, we also project them into the binary space in Line 12. After $T$ rounds of iterations, we craft adversarial malware samples for this batch in Line 13. (v) We repeat the above steps on each batch of the dataset to obtain enough diversified adversarial examples as candidates for adversarial purification.
@@ 4.3 Protective Noise Injection
The proposed diversified adversarial perturbation mechanism successfully enhances the diversity of adversarial malware samples within the training dataset, however, if benign samples are left unprocessed, the purification model may have a tendency to misclassify them as malicious samples, resulting in erroneous classification outcomes. It is also important to note that adversaries typically do not alter benign samples to make them appear as malware. As a result, the direct application of adversarial perturbation methods to manipulate benign samples is not viable.
To solve this problem and protect clean samples from excessive purification, we here propose a projective noise injection mechanism that preprocesses clean data before the training phase of the purification model. Specifically, we introduce a threshold $\eta\in[0,1]$ to control the range of noise injection. When $\eta$ is set to 0, no noise is added, while a value of 1 indicates that all feature values will be flipped under this condition. By employing this approach, the purification model is encouraged to better discriminate between benign and malicious samples within the feature space, thereby ensuring the detection performance of subsequent classifiers on clean data.
The detailed algorithm of the protective noise injection is described in Algorithm 2. (i) For each batch of training dataset, we first select benign samples $(x_{i},y_{i}=0)$ in Line 2, where $i\in[1,N]$. (ii) Then, we calculate the batch size $b_{i}$ and length $l_{i}$ of this batch of benign data for the following noise injection. (iii) We randomly generate a mask and determine the injection position according to the preset noise level $\eta$. (iv) We modify binary feature values on these positions (flipping from ”0” to ”1” and from ”1” to ”0”) based on these generated masks. (v) Finally, by applying the aforementioned steps to all batches of the training dataset, we obtain a subset of benign samples that have been augmented with protective noise.

### v2

@@ Abstract
Machine learning (ML) has gained significant adoption in Android malware detection to address the escalating threats posed by the rapid proliferation of malware attacks. However, recent studies have revealed the inherent vulnerabilities of ML-based detection systems to evasion attacks. While efforts have been made to address this critical issue, many of the existing defensive methods encounter challenges such as lower effectiveness or reduced generalization capabilities. In this paper, we introduce MalPurifier, a novel adversarial purification framework specifically engineered for Android malware detection. Specifically, MalPurifier integrates three key innovations: a diversified adversarial perturbation mechanism for robustness and generalizability, a protective noise injection strategy for benign data integrity, and a Denoising AutoEncoder (DAE) with a dual-objective loss for accurate purification and classification. Extensive experiments on two large-scale datasets demonstrate that MalPurifier significantly outperforms state-of-the-art defenses. It robustly defends against a comprehensive set of 37 perturbation-based evasion attacks, consistently achieving robust accuracies above 90.91%. As a lightweight, model-agnostic, and plug-and-play module, MalPurifier offers a practical and effective solution to bolster the security of ML-based Android malware detectors.
@@ 1 Introduction
Despite these advances in detection capabilities, ML-based detectors are vulnerable to adversarial examples, which are created by modifying non-functional instructions in executable programs of existing malware [8, 9]. Adversarial examples can enable a range of attacks, including evasion attacks [10, 11, 12], poisoning attacks [13, 14, 15], or a combination of both [16]. In this study, we specifically narrow our focus to evasion attacks, which are designed to deceive ML-based detection during the testing phase.
@@
Another defense technique, known as adversarial purification [23, 24, 25], aims to remove potential perturbations from input samples, resulting in purified samples that can be correctly classified by the target classifier. The purification model is usually trained independently of the classification model and does not necessarily require class labels [26]. As a result, it can mitigate unseen threats in a plug-and-play manner without re-training the target classifier [27], leading to less training overhead and more flexible employment. Nevertheless, existing purification solutions for image classification are not easily applicable to Android malware detection due to significant differences [28] as: (i) The feature space of Android applications is not only discrete but also high-dimensional, making noise removal fundamentally different from denoising continuous pixel values in the image domain. (ii) Evasion attacks in this domain are far more diverse than simple perturbations, including complex structural and semantic manipulations that require more than a simple reconstruction objective to defend against. (iii) A practical defense must maintain high accuracy on benign samples, as false positives can render security products unusable. However, this constraint is less stringent in other domains.
To overcome these challenges, we propose MalPurifier, a novel adversarial purification framework specifically engineered for the Android malware ecosystem, as illustrated in Fig. 1. MalPurifier introduces three key innovations to handle the diverse and complex attack vectors: (i) To defend against a wide spectrum of unforeseen attacks, we introduce a mechanism that generates adversarial malware with progressively increasing perturbation strengths. This forces the purification model to learn a more generalizable representation of maliciousness, rather than overfitting to a specific attack type. (ii) To prevent the degradation of accuracy on clean data, we propose a novel strategy that injects “protective noises” into benign samples. This teaches the purifier to preserve the integrity of legitimate applications and avoid costly false positives. (iii) We design a Denoising AutoEncoder (DAE) with a tailored loss function that uniquely combines both reconstruction and feature-space prediction objectives. This dual-objective approach ensures that purified samples are not only restored to their original form but are also correctly interpreted by the downstream detector. As a result, MalPurifier operates as a non-intrusive, computationally efficient, and easily scalable plug-and-play module that significantly enhances the robustness of existing malware detectors. The main contributions of this paper can be summarized as follows:
  • Novel purification framework for Android evasion attacks. We propose MalPurifier, a robust purification framework tailored for the unique challenges of the Android ecosystem. Instead of a generic application of autoencoders, our method introduces novel mechanisms specifically designed to handle the discrete feature space and diverse attack vectors inherent to Android malware.
@@
  • Trade-off between robustness and detection accuracy. Our proposed mechanism injects a diversified range of perturbations into malware samples, ranging from no injection to worst-case perturbations, to enhance robustness against different attacks, especially unforeseen attacks. We also implement a novel protective noise injection strategy to prevent false positives on benign samples.
@@
  • Experimental validation on public datasets. We compare MalPurifier with the state-of-the-art (SOTA) methods via Drebin [29] and Androzoo [30] datasets. Experimental results show that MalPurifier significantly outperforms other defenses against a comprehensive set of perturbation-based and structure-based evasion attacks, with less overhead required. Finally, we release the source code at https://github.com/SEU-ProactiveSecurity-Group/MalPurifier.
@@ 2 Preliminaries
This section provides the necessary background for understanding our approach. We first review ML-based Android malware detection in Section 2.1, then examine various evasion attack methods in Section 2.2, and finally introduce the concept of adversarial purification in Section 2.3.
@@ 2.3 Adversarial Purification
To counter these diverse evasion attacks, adversarial purification [39] has emerged as a promising defense strategy. The fundamental concept behind it is to preprocess the input data directly, preventing any embedded adversarial components from feeding into the target model, so that the influence of attacks can be mitigated. These methods are widely regarded as model-agnostic and highly efficient, making them easy to train and utilize while demonstrating strong generalization capabilities.
@@ 3.1 Threat Model
To perform analysis on evasion attacks and develop countermeasures, we consider the threat model in terms of assumptions regarding the attacker’s capabilities and knowledge of the target system. Especially, we discuss attack scenarios specified by three different threat models in this section.
@@ 3.1.1 Black-Box Attacks
In this attack scenario, the attacker has no knowledge about either the malware detector $f$ or the purifier $g$ (e.g., obfuscation attacks). Here, the adversary has to rely on trial and error to find perturbations based on the limited feedback. Given a malware example $z$, the attacker attempts to perturb it from $z$ to $z^{\prime}$, resulting in a feature space transformation from $x=\phi(z)$ to $x^{\prime}=\phi(z^{\prime})$. The attacker’s goal is to make the prediction of the target system incorrect, such that
@@
where the range set of $[\check{u},\hat{u}]$ ensures the feasibility of manipulations and the persistence of malicious functions.
@@ 3.1.2 Grey-Box Attacks
In the case of grey-box attacks, since the adversary is aware of the characteristics of the original malware classifier $f$ but is oblivious to the presence and structure of the purifier $g$, we also refer to this attack pattern as oblivious attacks. Hence, attack strategy in this scenario focuses on deceiving the unsecured classifier $f$ using adversarial examples generated from malicious examples without considering the purifier $g$. Formally, given a malware example $z$ with its feature representation $x=\phi(z)$ and label $y=1$, the attacker will modify it to obtain the adversarial feature representation $x^{\prime}$ that can evade detection, by solving
@@ 3.1.3 White-Box Attacks
In the white-box attack setting, the attacker is granted complete knowledge of both the target model $f$ and the purifier’s architecture $g$. This means the adversary is fully aware of the entire defense mechanism and can dynamically adapt strategies according to it, thus constituting what is formally known as adaptive attacks. In this scenario, the adversary aims to craft more sophisticated attacks that mislead both the malware detector and adversarial purifier simultaneously. It is worth noting that the sample will be first feed into the purifier and its output will be then classified by the malware detector. Therefore, given a malware feature-label pair $(x,y)$, the attacker needs to perturb $x$ into $x^{\prime}$ by solving
@@ 3.2 Defense Formulation
  • High-accuracy Android malware detection. Developing a malware detection model that can classify clean samples into benign or malicious with high accuracy, that is, $\varphi_{\theta}(x)=y$, in which $x$ can be the feature representation of a normal Android application or a malware sample, and thus $y=0$ or $1$. This model can be easily obtained by utilizing some ML algorithms (e.g., Deep Neural Network (DNN)), which have shown promising results with 99% accuracy in their laboratory settings [40, 11, 6].
@@ 4 The MalPurifier Approach
In this section, we present the MalPurifier approach in detail. We first provide an overview of the system architecture, including both the training and inference processes. Next, we describe the mechanisms for generating diversified adversarial perturbations and injecting protective noise. These mechanisms are crucial for creating robust training data for the purification model. Finally, we introduce the label-independent DAE-based purification model, which is responsible for producing purified samples and improving detection accuracy.
@@
4.1 Architecture Overview
Figure 2 provides an overview of the MalPurifier architecture. As we can see, it is composed of three main modules: (i) a feature extractor $\phi(\cdot)$ that maps an Android application into a feature vector, (ii) an adversarial purifier $g$ that processes samples via a DAE model $\psi_{\vartheta}$, and (iii) a malware detector $f$ that uses a DNN model $\varphi_{\theta}$ for detection.
During the training phase, we first extract features from batches of clean (or natural) data. These samples include both benign and malicious examples without prior manipulation. These features are then fed into the detector model (e.g., DNN), which iteratively updates its parameters to minimize the loss function. Once the DNN model is effectively trained, it exhibits exceptional classification accuracy when presented with clean inputs. Notably, we do not retrain the DNN with labeled adversarial data. This key aspect distinguishes our approach from traditional adversarial training methods.
In contrast, MalPurifier incorporates a purification model (e.g., DAE). This model is specifically designed to learn compact representations of input data and reconstruct the original (clean) data from its noisy or perturbed versions. The effectiveness of the DAE largely depends on the quality and diversity of the training data, particularly the types and levels of noise introduced during training.
To address this, we first propose a diversified adversarial perturbation mechanism (see Section 4.2). This mechanism generates adversarial malware samples with varying degrees of perturbation, enabling the DAE to learn how to handle a broad spectrum of adversarial manipulations. As a result, the purification model is better equipped to mitigate different types of evasion attacks. Then, we propose a noise injection strategy for benign samples (see Section 4.3). This approach helps prevent the DAE from over-correcting or inadvertently altering clean samples, thereby preserving their original feature patterns.
In addition, we incorporate both reconstruction loss and prediction loss into the DAE’s objective function. The model is trained using a combination of adversarially perturbed malware samples, noisy benign samples, and their corresponding clean counterparts. Importantly, the DAE is trained independently of class labels, which enhances its generalization capability. Further details on the DAE training process are provided in Section 4.4.
During the testing phase, an input sample is sequentially processed by MalPurifier’s feature extraction, purification, and detection modules to yield a prediction result, as illustrated in the right part of Fig. 2.
@@ 4.2 Diversified Adversarial Perturbation
To address the challenge of diverse attack vectors in the malware domain, a simple perturbation strategy (e.g., adversarial training) is insufficient. To this end, we propose a diversified adversarial perturbation mechanism designed to enhance the generalizability of the purification model.
By exposing the model to a wide variety of perturbations, it enables the purifier to defend not only against known attack types but also to effectively mitigate previously unseen or unknown attacks. Therefore, our approach aims to maximize the difference in feature space between the original malware sample $x$ and its perturbed counterpart $x^{\prime}$. This objective can be formally expressed as
$\Delta(x,x^{\prime})=d(\mathcal{F}_{\theta}(x)|n,\mathcal{F}\theta(x^{\prime})|n),\ \ \mathrm{s.t.}x^{\prime}\in[\check{u},\hat{u}],$ (9)
where $\mathcal{F}\theta(x)|_{n}$ is the internal feature representation of $x$ at the $n$th layer of the malware detector $f$, and $d(\cdot)$ denotes the distance metric used to quantify the difference between the original and perturbed features. In our implementation, we use Mean Square Error (MSE) as the distance metric.
The process for generating diversified adversarial examples is outlined in Algorithm 1. The main steps are as follows: (i) We begin by sampling a batch of malware examples $(x_{i},y_{i}=1)_{i=1}^{N}$ in Line 2. (ii) For each batch, we set the adversarial depth in proportion to the batch index in Line 3, so that the perturbation level is gradually increased across batches to cover different attack intensities. (iii) For the first batch, no perturbations are applied in Line 5. (iv) For subsequent batches, we initialize $x_{i}$ with a random transformation in Line 7. We then compute the gradient $g_{t}$ based on the feature difference in Line 10 and iteratively update the sample according to the preset adversarial depth in Line 11. (v) To ensure practicality, each generated adversarial sample is projected back into the binary feature space in Line 12. (vi) After $T$ iterations, we obtain the final adversarial samples for the batch in Line 13. (vii) By repeating these steps for all batches, we obtain a comprehensive set of adversarial examples with varying perturbation strengths, ensuring both diversity and generalizability for robust purification.
@@ 4.3 Protective Noise Injection
While the diversified perturbation strategies described in Section 4.2 can significantly enhance the generalizability of the purification model against various evasion attacks, they may also introduce a new critical challenge, such that the risk of over-purification on clean data. Excessive purification can inadvertently corrupt benign samples and increase the false positive rate. Moreover, it is important to note that, in practice, adversaries rarely attempt to modify benign samples to mimic malware. Therefore, directly applying adversarial perturbations to benign data is both unrealistic and unnecessary.
To tackle this, we introduce a novel protective noise injection mechanism, a key contribution for ensuring the practical usability of the defense. Specifically, it introduces controlled random noise, enhancing the purification model’s ability to correctly process benign samples. Here, we define a threshold parameter $\eta\in[0,1]$ to control the extent of noise injection ($\eta=0$ means no noise added while $\eta=1$ indicates that all features are flipped). By tuning $\eta$, we can balance the trade-off between robustness and accuracy, ensuring that the purification model maintains high detection performance on clean data.
Algorithm 2 presents the detailed steps of our protective noise injection mechanism: (i) For each batch in the training dataset, we first select all benign samples $(x_{i},y_{i}=0)$ in Line 2. (ii) We then determine the batch size $b_{i}$ and the feature length $l_{i}$ for subsequent processing in Line 3. (iii) Next, we randomly generate a binary mask based on the preset noise level $\eta$ to identify which feature positions will be altered in Line 4. (iv) The selected feature values are then flipped (from ”0” to ”1” or vice versa) according to the generated mask in Line 5. (v) By repeating these steps for all batches, we obtain an augmented set of benign samples with protective noise in Line 6. These modified benign samples are subsequently used to train the purification model.